OpenStudy (lena772):

Two systems of equations are shown below: System A 3x + 2y = 3 –2x - 8y = –1 System B –x - 14y = 1 –2x – 8y = –1 Which of the following statements is correct about the two systems of equations? The value of x for System B will be one–third of the value of x for System A because the coefficient of x in the first equation of System B is one third times the coefficient of x in the first equation of System A.

OpenStudy (lena772):

The value of x for System A will be equal to the value of y for System B because the first equation of System B is obtained by adding –4 to the first equation of System A and the second equations are identical. They will have the same solution because they represent the same lines when plotted on the coordinate axes. They will have the same solution because the first equation of System B is obtained by adding the first equation of System A to 2 times the second equation of System A.

OpenStudy (lena772):

Maybe you multiply an equation to eliminate the y?

OpenStudy (lena772):

4(3x + 2y = 3) –2x - 8y = –1 ??

OpenStudy (ranga):

multiply it out in the first equation and add the 2 equation.

OpenStudy (lena772):

12x+8y=12 -2x-8y=-1 10x=11 x=1.1

OpenStudy (ranga):

yes. might as well solve the whole thing. find y.

OpenStudy (lena772):

13.2+8y=12 8y=-1.2 y=-0.15 ?

OpenStudy (ranga):

yes. And do the same for system B. solve for x and y.

OpenStudy (lena772):

2(–x - 14y = 1) –2x – 8y = –1

OpenStudy (ranga):

I don't think I read the choices correctly. But we will soon find out.

OpenStudy (lena772):

-2x-28y=2 -2x-8y=-1 -20y=3 y=-0.15

OpenStudy (ranga):

yes. I did not read the choices correctly. Check out the last choice.

OpenStudy (lena772):

-x-2.1=1 -x=3.1 x=-3.1

OpenStudy (lena772):

But i got a different x value for b

OpenStudy (ranga):

the x is not right.

OpenStudy (lena772):

-x+2.1=1 -x=-1.1 x=1.1 ?

OpenStudy (lena772):

So it's D?

OpenStudy (ranga):

yes.